The partition coefficient of a base B (monoacidic) at a particular pH, in an octanol-water mixture is 1.45. Consider an experiment where 3.5 g of B are shaken with 100 mL of octanol and 50 mL of water.
Molar mass of B = 101.19 g mol-1

(i) Calculate the mass of B in the aqueous layer after shaking.
1
Expert's answer
2020-06-08T15:49:33-0400

nB = 3.5/101.19 = 0.035 mol ; nin octaniol = x mol; nin water = 0.035 - x mol

Cin water = (0.035 - x)/0.050 mol/L

Cin octanol = x/0.100 mol/L

KD = Cin octanol/Cin water

1.45 = x*0.050/((0.035 - x)*0.100)

x = 0.026 mol

nin water = 0.035 - 0.026 = 0.0090 mol

min water = 0.0090*101.19 = 0.91 g